OpenStudy (anonymous):

Help please! How may permutations of the word “spell” are there? There are ________ways.

OpenStudy (misty1212):

lol hi again!

OpenStudy (misty1212):

how many letters are there?

OpenStudy (misty1212):

that is not a hard question, just count them, then we can go from there

OpenStudy (anonymous):

5

OpenStudy (misty1212):

ok good so if all the letters were different (they are not) then the answer would be \[5!=5\times 4\times 3\times 2\]

OpenStudy (misty1212):

but you have two letters that are the same, namely the "l"

OpenStudy (misty1212):

so in this case it is \[\frac{5!}{2}=5\times 4\times 3\](the 2's cancel)

OpenStudy (anonymous):

which the answer is??

OpenStudy (misty1212):

whatever you get when you multiply

OpenStudy (anonymous):

i got 60 correct?

OpenStudy (anonymous):

i got 60 too?

OpenStudy (anonymous):

60

OpenStudy (anonymous):

i swear its 60

OpenStudy (anonymous):

i know it is

OpenStudy (misty1212):

lol yeah \[\huge\color{red}{60}\]

OpenStudy (anonymous):

misty right? <3

OpenStudy (anonymous):

Thanks Misty you're awesome

OpenStudy (anonymous):

your a hero

OpenStudy (anonymous):

lol